OpenStudy (anonymous):

Solve the equation cos2x = 1-sinx for x in the interval [0,2pi). I used the double angle formula to get cos^2x - sin^2x= 1- sin x then i moved it all to the left side and set it equal to 0: cos^2x - sin^2x + sinx - 1= 0. Have I done this correctly so far? Also where do I go next?

OpenStudy (anonymous):

ok from there you can use the identity cos^2x = 1 - sin^2x

OpenStudy (anonymous):

so the equation will now be 1 - sin^2x - sin^2x + sinx - 1

OpenStudy (anonymous):

then do i combine like terms?

OpenStudy (anonymous):

yes you do and you should get a quadratic equation

OpenStudy (anonymous):

do i get -2sin^2 + sin =0

OpenStudy (anonymous):

yes and now you factorise

OpenStudy (anonymous):

then do i get sinx(-2sinx+1)=0

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now sinx = 0 and -2 sinx +1=0 and you can solve from there

OpenStudy (anonymous):

yupp i did that and do i get 0, pi, pi/6 and 5pi/6

OpenStudy (anonymous):

i think you should just get 0 and pi/3 because -2 sinx +1 =0 then -2sinx=-1 then sinx= -1/-2 which will be sinx=1/2 which is 30degrees pi/3

OpenStudy (anonymous):

oh you are right for some reason i divided it by 2 instead of -2 thank you!!!!

OpenStudy (anonymous):

pi/6 sorry

OpenStudy (anonymous):

wait but its a multiple choice question and the answers are: 0, pi pi/6, 5pi/6 0, pi, 7pi/6, 11pi/6 0, pi, pi/6, 5pi/6

OpenStudy (anonymous):

wait i think im right with the answers. because you get sin=0 which is 0, pi and you get sin=1/2 which is pi/6, 5pi/6

OpenStudy (anonymous):

yes you are

OpenStudy (anonymous):

okay thank you for all your help!

OpenStudy (anonymous):

no problem